OpenStudy (anonymous):

8. Which of the following is the equation of the line passing through points A(−5, 2) and B(0, −8)? (1 point) a. y = −2x + 8 / b.y = 2x + 8/ c. y = −2x − 8/ d. y = 2x − 8

OpenStudy (anonymous):

8. Which of the following is the equation of the line passing through points A(−5, 2) and B(0, −8)? (1 point) a. y = −2x + 8 b.y = 2x + 8 c. y = −2x − 8 d. y = 2x − 8